# SymCERE: Symmetric Contrastive Learning for Robust Review-Enhanced Recommendation

### v2

@@ 6. Discussion
Fundamentally, the triad of challenges we identified—false negatives, popularity bias, and signal ambiguity—are not just concurrent but deeply intertwined. Popularity bias, by encoding popularity as vector magnitude, can severely distort the embedding space. This distortion exacerbates the false negative problem, as true semantic similarity (based on vector direction) is masked by the dominant, non-semantic popularity signal. A model struggling with this skewed space is then ill-equipped to resolve signal ambiguity; it may learn to rely on the strong but noisy popularity signal rather than discerning subtle, objective cues from text. This cascade effect underscores why a unified approach is critical. Addressing these issues in isolation is insufficient, as the unresolved problems will continue to undermine the effectiveness of partial solutions. Our work, therefore, provides a holistic solution by structurally tackling these interconnected challenges simultaneously.
@@ 6.1. A Geometric View of Debiasing and Performance Gains (RQ2)
First, we posited a ”magnitude-as-noise” hypothesis. Our ablation studies (Table 3) show that a model variant without L2 normalisation performs poorly. This is particularly evident with the non-linear NGCF backbone, where performance drops to near-zero levels. A likely explanation for this catastrophic failure is that without the constraint of normalisation, the embedding magnitudes can grow uncontrollably during training due to NGCF’s non-linear transformations, leading to numerical instability and model divergence. This underscores the critical role of L2 normalisation in maintaining a stable training process, especially for more complex GNN architectures. In contrast, this is likely because, without normalisation, the L2-norm of popular items can dominate similarity calculations, potentially masking the directional preference signal.
@@
L2 normalisation addresses this issue and provides the foundation for improving the effectiveness of contrastive learning. We theorized that normalisation improves Uniformity by spreading embeddings more evenly across the unit hypersphere to maximize representational capacity. This claim is empirically supported by the data in Table 4. Across all datasets, applying L2 normalisation increased the standard deviation of cosine similarities, providing quantitative evidence that the embeddings are more uniformly distributed, thus forming a more discriminative representation space.
These geometric improvements—enhanced uniformity and dimensional utilization—are mechanisms that can suppress noise and bias, leading to gains in recommendation performance.
Interestingly, our ablation study (Table 3) also reveals that the InfoNCE-based model variant occasionally outperforms the SINCERE-based model when using the non-linear NGCF backbone. We hypothesize this is a result of implicit regularization; while InfoNCE’s treatment of false negatives is flawed, the resulting noisy gradient may prevent a complex model like NGCF from overfitting. In contrast, SINCERE’s theoretically cleaner signal might not provide the same regularizing effect on certain datasets. This suggests a nuanced trade-off between the theoretical correctness of a loss function and its practical behavior with different model architectures.
@@ 6.2.3. Low-Improvement Case: Clothing, Shoes and Jewelry
Subjective Aesthetics ‘beautiful‘, ‘elegant‘, ‘style‘, ‘looks‘ Highly Ambiguous Personal taste, no objective benchmark
Context-Dependent Fit ‘size‘, ‘fit‘, ‘small‘, ‘large‘ Ambiguous Depends on user’s body/preference
@@ 8. Conclusion
We addressed the interconnected challenges of false negatives, popularity bias, and signal ambiguity with SymCERE, a unified contrastive learning method. By integrating a symmetric SINCERE loss with hyperspherical projection (L2 normalisation), SymCERE significantly outperforms strong baselines on both NDCG@10 and HR@10 across 15 datasets (e.g., +43.6% NDCG@10). Our key finding, termed ‘semantic anchoring’, reveals the model learns to prioritize objective, information-rich, and domain-specific attributes over superficial sentiment, showing that robust alignment stems from factual understanding. This unified and principled approach of structurally mitigating biases thus creates a path toward recommender systems that are not only more accurate but also more robust, interpretable, and fair by promoting item diversity.

### v3

@@ 6.1. A Geometric View of Debiasing and Performance Gains (RQ2)
First, we posited a ”magnitude-as-noise” hypothesis. Our ablation studies (Table 3) show that a model variant without L2 normalisation performs poorly. This is particularly evident with the non-linear NGCF backbone, where performance drops to near-zero levels. This confirms that without spherical projection, popularity bias (encoded as magnitude) dominates the embedding space, drowning out the semantic directional signals. L2 normalization is thus a structural prerequisite.
@@
L2 normalisation addresses this issue and provides the foundation for improving the effectiveness of contrastive learning. We theorized that normalisation improves Uniformity by spreading embeddings more evenly across the unit hypersphere to maximize representational capacity. This claim is empirically supported by the data in Table 4 and visualized in Figure 4. Across all datasets, applying L2 normalisation increased the standard deviation of cosine similarities and the variance of dimensions, providing quantitative and visual evidence that the embeddings are more uniformly distributed, thus forming a more discriminative representation space.
Figure 4. Impact of L2 Normalization on Dimensional Utilization (Pet Supplies). The orange distribution (With Normalization) exhibits significantly higher variance across dimensions compared to the blue distribution (Without Normalization). This confirms that SymCERE prevents dimensional collapse, utilizing the embedding space more uniformly to encode semantic differences.
Table 3. Analysis of Fusion Mechanisms (Ablation Study). $M$ w/ InfoNCE represents the standard CL fusion baseline (i.e., without false negative removal). w/o Norm removes L2 normalization. ”Min. Drop” indicates the percentage drop in performance relative to the full model (SymCERE w/ Norm).
Dataset Metric NGCF Backbone LightGCN Backbone
SymCERE
w/o Norm
InfoNCE
w/ Norm
SymCERE
w/ Norm
Min. Drop
SymCERE
w/o Norm
InfoNCE
w/ Norm
SymCERE
w/ Norm
Min. Drop
Clothing, Shoes & Jewelry HR@10 0.002 0.536 0.538 $\downarrow 99.6\%$ 0.548 0.547 0.550 $\downarrow 0.6\%$
NDCG@10 0.001 0.428 0.425 $\downarrow 99.8\%$ 0.385 0.384 0.398 $\downarrow 3.6\%$
CDs & Vinyl HR@10 0.007 0.490 0.488 $\downarrow 98.6\%$ 0.492 0.463 0.495 $\downarrow 6.4\%$
NDCG@10 0.002 0.338 0.332 $\downarrow 99.4\%$ 0.303 0.295 0.308 $\downarrow 4.2\%$
Tools & Home Imp. HR@10 0.005 0.397 0.402 $\downarrow 98.8\%$ 0.409 0.396 0.412 $\downarrow 3.8\%$
NDCG@10 0.003 0.186 0.187 $\downarrow 98.4\%$ 0.156 0.153 0.162 $\downarrow 6.0\%$
Sports & Outdoors HR@10 0.002 0.426 0.424 $\downarrow 99.5\%$ 0.433 0.422 0.434 $\downarrow 2.9\%$
NDCG@10 0.001 0.227 0.226 $\downarrow 99.6\%$ 0.187 0.184 0.195 $\downarrow 5.6\%$
Beauty & Personal Care HR@10 0.001 0.390 0.386 $\downarrow 99.7\%$ 0.401 0.374 0.405 $\downarrow 7.7\%$
NDCG@10 0.000 0.158 0.162 $\downarrow 100\%$ 0.128 0.122 0.135 $\downarrow 9.6\%$
Automotive HR@10 0.005 0.450 0.446 $\downarrow 98.9\%$ 0.469 0.459 0.470 $\downarrow 2.3\%$
NDCG@10 0.002 0.283 0.285 $\downarrow 99.3\%$ 0.254 0.252 0.265 $\downarrow 4.9\%$
Toys & Games HR@10 0.001 0.417 0.415 $\downarrow 99.8\%$ 0.416 0.407 0.419 $\downarrow 3.0\%$
NDCG@10 0.000 0.197 0.198 $\downarrow 100\%$ 0.164 0.162 0.173 $\downarrow 6.3\%$
Pet Supplies HR@10 0.004 0.371 0.373 $\downarrow 98.9\%$ 0.383 0.369 0.387 $\downarrow 4.6\%$
NDCG@10 0.001 0.124 0.129 $\downarrow 99.2\%$ 0.096 0.093 0.102 $\downarrow 8.9\%$
Health and Household HR@10 0.002 0.355 0.358 $\downarrow 99.4\%$ 0.343 0.291 0.357 $\downarrow 18.5\%$
NDCG@10 0.001 0.130 0.131 $\downarrow 99.2\%$ 0.102 0.098 0.112 $\downarrow 12.5\%$
Video Games HR@10 0.006 0.331 0.332 $\downarrow 98.2\%$ 0.280 0.222 0.326 $\downarrow 31.9\%$
NDCG@10 0.002 0.129 0.130 $\downarrow 98.5\%$ 0.099 0.089 0.116 $\downarrow 23.3\%$
Cell Phones and Acc. HR@10 0.005 0.397 0.408 $\downarrow 98.8\%$ 0.408 0.357 0.401 $\downarrow 11.0\%$
NDCG@10 0.001 0.170 0.178 $\downarrow 99.4\%$ 0.146 0.149 0.161 $\downarrow 7.5\%$
Patio, Lawn and Garden HR@10 0.007 0.440 0.433 $\downarrow 98.4\%$ 0.440 0.447 0.450 $\downarrow 2.2\%$
NDCG@10 0.002 0.217 0.225 $\downarrow 99.1\%$ 0.179 0.182 0.189 $\downarrow 5.3\%$
Office Products HR@10 0.004 0.430 0.430 $\downarrow 99.1\%$ 0.430 0.422 0.430 $\downarrow 1.8\%$
NDCG@10 0.002 0.203 0.190 $\downarrow 98.9\%$ 0.147 0.147 0.161 $\downarrow 8.8\%$
Yelp HR@10 0.017 0.350 0.344 $\downarrow 95.1\%$ 0.224 0.210 0.293 $\downarrow 28.3\%$
NDCG@10 0.003 0.113 0.112 $\downarrow 97.3\%$ 0.070 0.068 0.089 $\downarrow 23.6\%$
Travel (JP) HR@10 0.040 0.789 0.768 $\downarrow 94.8\%$ 0.829 0.695 0.824 $\downarrow 15.7\%$
NDCG@10 0.006 0.298 0.269 $\downarrow 97.8\%$ 0.390 0.228 0.380 $\downarrow\textbf{40.0\%}$
@@ 6.2.3. Low-Improvement Case: Clothing, Shoes and Jewelry
Product Spec/Function ‘size‘, ‘fit‘, ‘small‘, ‘large‘ Ambiguous Depends on user’s body/preference
Subjective Experience ‘beautiful‘, ‘elegant‘, ‘style‘, ‘looks‘ Highly Ambiguous Personal taste, no objective benchmark
@@ 8. Conclusion
We addressed the interconnected challenges of false negatives, popularity bias, and signal ambiguity with SymCERE, a unified contrastive learning method. By integrating a symmetric NCE loss with hyperspherical projection (L2 normalisation), SymCERE significantly outperforms strong baselines on both NDCG@10 and HR@10 across 15 datasets (e.g., +43.6% NDCG@10). Our key finding, termed ‘semantic anchoring’, reveals the model learns to prioritize objective, information-rich, and domain-specific attributes over superficial sentiment, showing that robust alignment stems from factual understanding. This unified and principled approach of structurally mitigating biases thus creates a path toward recommender systems that are not only more accurate but also more robust, interpretable, and fair by promoting item diversity.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP23K28908, JP25K21210.